jbbenedict:

The cost of hiring a catering service to serve for a party is 150php per head for 20 persons or less. 130php per head for 21 to 50 persons, and 110php per head for 51 to 100 persons, for 100 or more the cost is 100php. Represent the total cost as a piecewise function of the number of atendees of the party.

ArtisticBeeTeen:

What do you mean php and per head?

dukedadon:

alternatively, it could also be written as: f(x) = 0 for x = 0 f(x) = 150 * x for 0 < x <= 20 f(x) = 130 * x for 20 < x <= 50 f(x) = 110 * x for 50 < x <= 100 f(x) = 100 * x for 100 < x <= infinity there are various ways to express it, but the main idea is that the function f(x) is defined differently for different intervals of x. let f(x) represent the cost function. let x represent the number of persons.

jbbenedict:

php means philippines peso and per head means the average per person

ArtisticBeeTeen:

oh ok...well nvm ig someone else HAD to help

dukedadon:

@jbbenedict wrote:
php means philippines peso and per head means the average per person
150 pesos per head for 20 persons or less. 130 pesos for 21 to 50 persons. 110 pesos per head for 51 to 100 persons. 100 pesos per head for 101 or more pesos.

dukedadon:

@artisticbeeteen wrote:
oh ok...well nvm ig someone else HAD to help
I gotchu man, continue your day (:

dukedadon:

your piece-wise function would look like: f(x) = 150 * x for 0 <= x <= 20 f(x) = 130 * x for 21 <= x <= 50 f(x) = 110 * x for 51 <= x <= 100 f(x) = 100 * x for x >= 101

jbbenedict:

thank you man!

dukedadon:

@jbbenedict wrote:
thank you man!
np :D

jbbenedict:

@dukedadon wrote:
@jbbenedict wrote:
thank you man!
np :D
F(x) = 150 if 1≥x≥20 F(x) = 130 if 21≥x≥50 F(x) = 110 if 51≥x≥100 F(x) = 100 if x>100

jbbenedict:

i think this is a answer?

dukedadon:

what is your final answer?

jbbenedict:

F(x) = 150 if 0<x<20 F(x) = 130 if 21<x<50 F(x) = 110 if 51<x,<100 F(x) = 100 if x>100